Chapter 06.01 Exercise Question 10 6.1.10 A sample of heights (in inches) of 20 female statistics students is as follows: 73, 62, 68, 64, 68, 64, 72,61, 59, 60, 70, 69, 79, 67, 67, 69, 63, 62, 63, 69 Find the median height. Round your answer to 1 decimal place, e.g. o.5 the absolute tolerance is +/-o Find the first and third quartiles. Round your answers to 1 decimal place, e.g. 0.5. , Q3 Find the inter-quartile range. Round your answer to 1 decimal places, e.g. o.5. IQR the absolute tolerance is +/-0.1
media%2Fb77%2Fb77167a6-1099-4498-a744-06
Answer #1

putting above data in ascending order:

S. no x
1 59
2 60
3 61
4 62
5 62
6 63
7 63
8 64
9 64
10 67
11 67
12 68
13 68
14 69
15 69
16 69
17 70
18 72
19 73
20 79

median height =middle value =(20+1)/2 th value =10.5 th value =(67+67)/2 =67

Q1 =npth value =20*0.25 th value =5th value =(62+63)/2 =62.5

Q3 =npth value =20*0.75 th value =15th value =(69+69)/2 =69

IQR =Q3-Q1 =69-62.5 =6.5

interquartile range =the length of the box